OpenStudy (anonymous):

This one is even harder! ): Given F(x)= 4x^3 and G(x)=2x what is F(x) divided by G(x)

OpenStudy (anonymous):

4x^3/(2x)= (4/2) * x^3/x= 2 * x^2

OpenStudy (anonymous):

oh okay so its 2x^2? @katherine.ok

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

thank you for your help !